# Observer Design over Hypercomplex Quaternions

### v2

@@ Abstract
We develop observer design over hypercomplex quaternions in a characteristic-polynomial-free framework. Using the standard right-module convention, we derive a right observable companion form and its companion polynomial that encodes error dynamics via right-eigenvalue similarity classes. The design mirrors the real/complex case—coefficient updates in companion coordinates, followed by a similarity back—yet avoids determinants, characteristic/minimal polynomials, and Cayley–Hamilton identities that do not transfer to quaternions. We also give an Ackermann-type construction for the important case of closed-loop companion polynomials with real coefficients, ensuring similarity-equivariant evaluation. The results yield simple recipes for full-order observers directly over quaternions, clarify the role of right spectra and their similarity classes, and pinpoint when classical one-shot formulas remain valid. Numerical examples illustrate the method and advantages over vectorized or complex-adjoint surrogates.
@@ I INTRODUCTION
In this paper, we develop an observer-focused, determinant–free framework built on observability and its duality with controllability. We construct the observable companion (observability-form) realization directly from the observability matrix, link its right spectrum to companion-polynomial zeros, and derive a quaternionic Observer Ackermann formula that assigns right–eigenvalue similarity classes using a desired real-coefficient polynomial. Throughout, we remain in the hypercomplex (scalar) model of $\mathbb{H}$ where signals, states, and operators live directly in $\mathbb{H}.$ We work with the right spectra, and use left substitution for the companion polynomial—choices that make noncommutativity explicit and usable.
In aerospace and robotics applications, by contrast, quaternions are typically treated in their real $4$-vector (scalar–vector) form, and observers/filters are implemented with purely real-vector operations and models [13]. These methods are highly effective for attitude estimation and remain the default in those domains.
Here, however, we adopt the hypercomplex definition $q\in\mathbb{H}$ and work directly with quaternionic states and operators. This preserves the noncommutative algebra essential for right/left spectra and similarity classes, for companion/observability forms, and for a central (real-coefficient) polynomial calculus $a(A)$ without coordinate artifacts. The viewpoint is also motivated by areas where the hypercomplex structure is intrinsic—e.g., quantum dynamics and control [5, 12], NMR/MRI pulse design [6], polarization/bivariate signal processing and color imaging [10], quaternion-valued filters [20], quantum computing [14] and emerging quaternionic learning models [19, 15]. We believe the observer results established here over $\mathbb{H}$—including the observable companion construction and the quaternionic Observer Ackermann formula—are original.
@@ III QUATERNIONIC STATE–SPACE SISO SYSTEMS
With $u\!\equiv\!0$, stability means $x(t)\!\to\!0$ (CT) or $x(k)\!\to\!0$ (DT) for all initial states. Equivalently, using the standard right eigenvalues $\lambda$ of $A$, the system is stable iff $\operatorname{Re}\lambda<0$ for all $\lambda$; (see [16]).
@@ IV OBSERVABLE COMPANION FORM AND COMPANION POLYNOMIAL
    Hence, the right spectrum of $A_{o}$ is exactly the union of the similarity classes of the right zeros of $a(\lambda)$:
    $\sigma_{\mathscr{r}}(A_{o})=\mathbin{{\cup}}_{i}[\lambda_{i}],$
    counting multiplicities.
@@
    Hence, the right spectrum of $A$ is exactly the union of the similarity classes of the right zeros of $a(\lambda)$: $\sigma_{\mathscr{r}}(A)=\mathbin{{\cup}}_{i}[\lambda_{i}]$, counting multiplicities.
@@
When $A$ is cyclic, it is natural to speak of the companion polynomial of $A$ constructed from any observable SISO pair $(A,C)$. Different cyclic choices of $C$ may yield different polynomials. Still, their right zeros agree up to quaternionic similarity classes: for every such $C$, the right zeros of the associated polynomial form the same multiset of classes—the right spectrum of $A$. For a fixed pair $(A,C)$, however, the observable companion $(A_{o},C_{o})$ and the attached polynomial are unique.
@@ V OBSERVERS
An observer is a dynamic estimator that reconstructs the state $\hat{x}$ from known inputs $u$ and measured outputs $y$ by injecting the innovation $y-C\hat{x}$ through a gain $L$. It enables implementable state feedback $u=K\hat{x}$, monitoring, and prediction. The relations below are standard from the real/complex literature and will be referenced later. They also hold for quaternionic systems over $\mathbb{H}$ with minor modifications that are detailed in a later section.
@@ VI DUALITY OVER $\mathbb{H}$
Over $\mathbb{R}$, transpose drives controller–observer duality: controllability of $(A,B)$ is equivalent to observability of $(A^{\mathsf{T}},B^{\mathsf{T}})$, and one often maps gains via $L=K^{\mathsf{T}}$ (cf. [2]).
Over $\mathbb{H}$, this fails because the correct adjoint is the conjugate transpose $M^{*}=\overline{M}^{\mathsf{T}}$, which alone preserves the quaternionic Hermitian form and converts left to right eigenpairs with eigenvalue conjugation [21, 18, 7, 1]. Thus, duality must be formulated with and with explicit care for right/left eigenpairs (see also quaternion-valued system treatments [9, 4]).
The original system and its dual system are related through mappings
$A\leftrightarrow A^{*},B\leftrightarrow C^{*},C\leftrightarrow B^{*},D\leftrightarrow D^{*}.$
Their controllability and observability are related via
@@
see [21, 18], and therefore
@@
If $K\in\mathbb{H}^{1\times n}$ is a controller for the dual pair $(A^{*},C^{*})$, then the observer (output-injection) gain for the original system is
@@
The error and dual closed-loop matrices satisfy $\sigma_{\mathscr{r}}\big(A-LC\big)=\sigma_{\mathscr{r}}\big(A^{*}-C^{*}K\big).$ Since $q]=[\overline{q}]$ for quaternionic similarity classes, the two spectra coincide as class sets. See [7, 1].
Proof sketches and references
The transpose fails over $\mathbb{H}$; only the adjoint satisfies $\langle Ax,y\rangle=\langle x,A^{*}y\rangle$ and aligns with quaternionic spectral theory [21, 18, 7, 1]. The dual mapping $(A,B,C,D)\!\leftrightarrow\!(A^{*},C^{*},B^{*},D^{*})$ is standard. Using $(AB)^{*}=B^{*}A^{*}$ and rank invariance yields $\mathcal{O}_{n}(A^{*},B^{*})=\mathcal{C}_{n}(A,B)^{*}$, hence $(A,B)$ controllable $\Leftrightarrow$ $(A^{*},B^{*})$ observable and $(A,C)$ observable $\Leftrightarrow$ $(A^{*},C^{*})$ controllable [21, 18]. Design $K$ on $(A^{*},C^{*})$ and set $L=K^{*}$; then $\sigma_{\mathscr{r}}(A-LC)=\overline{\sigma_{\mathscr{r}}(A^{*}-C^{*}K)}=\sigma_{\mathscr{r}}(A^{*}-C^{*}K)$ by conjugation invariance of right-eigenvalue classes [7, 1].
@@ VII QUATERNIONIC OBSERVER POLE PLACEMENT VIA COMPANION FORM
$A_{\mathrm{obs}}$ $:=$ $A-LC$ (24)
equals the multiset of right zeros (similarity classes, with multiplicities) of $a_{d}$. Moreover, with $L_{o}:=T^{-1}L$ we have
@@
has the finite right spectrum $\sigma_{\mathscr{r}}(A_{\rm obs})\!=\!\{-1,\!-2\}$ as desired.
@@
Here $\{-1\!+\!\mathbf{i},-1\!-\!\mathbf{i}\}\subset\sigma_{\mathscr{r}}(A_{\mathrm{obs}})$; equivalently, the entire spherical (similarity) class $\left[-1+\mathbf{i}\right]$ is assigned as expected.
(c) Quaternionic coefficients: Demanding specific quaternion poles (e.g., $\lambda_{1}=-1+\mathbf{j}$ and $\lambda_{2}=-2+\mathbf{k}$) sets (rounded)
@@
Its right spectrum is $\sigma_{\mathscr{r}}(A_{\mathrm{obs}})=\left[-1+\mathbf{i}\right]\cup[-2+\mathbf{i}]=\left[-1+\mathbf{j}\right]\cup\left[-2+\mathbf{k}\right]$ so $1\!+\!\mathbf{j},-2\!+\!\mathbf{k}\in\sigma_{\mathscr{r}}(A_{\mathrm{obs}})$ confirming that the observable-form coefficient update still assigns the desired similarity classes even when the coefficients are quaternionic.
@@ VIII OBSERVER GAIN VIA ACKERMANN’S FORMULA
Having established the quaternionic observer in its canonical (observable) form, we now turn to a compact algebraic method for directly computing the observer gain. The resulting Observer Ackermann’s formula is the dual counterpart of the control-side expression. It provides a one–shot computation of (L) achieving any admissible placement of right–eigenvalue classes (observer poles). Its structure mirrors the classical real/complex case. Yet, its justification over $\mathbb{H}$ requires special care: the argument avoids determinants and adjugates and instead exploits the observable companion relations and the right–polynomial calculus developed earlier. This formulation establishes the quaternionic duality between controllability and observability in a determinant–free setting.
@@
Let $A\in\mathbb{H}^{n\times n}$, $C\in\mathbb{H}^{1\times n}$ be such that the SISO pair $(A,C)$ is observable and the observability matrix $\mathcal{O}$ from (3) is invertible. Let $a_{d}(\lambda)=d_{0}+d_{1}\lambda+\cdots+d_{n-1}\lambda^{n-1}+\lambda^{n}\in\mathbb{R}[\lambda]$ be the desired monic error-dynamics polynomial with real coefficients, and let $e_{n}=[0,0,\dots,1]^{\mathsf{T}}\in\mathbb{R}^{n}$. Then the observer gain column
@@
$\sigma_{\mathscr{r}}(A-LC)$ $=$ $\mathbin{{\cup}}_{i=1}^{n}[\lambda_{i}],$ (29)
where $\{\lambda_{i}\}_{i=1}^{n}$ are the right roots of $a_{d}$ (listed with multiplicity).
@@
The restriction $a_{d}\in\mathbb{R}[\lambda]$ is essential: real coefficients lie in the center $Z(\mathbb{H})=\mathbb{R}$, so $a_{d}(A)$ is unambiguous and the required polynomial–similarity interchange is valid. The assigned classes are real or complex-conjugate spherical: for an $n$-th order SISO design, each real class consumes one degree and each nonreal class two, so $2s+r\leq n$ when placing $s$ nonreal and $r$ real classes.
@@
Similarity preserves the right spectrum, hence $\sigma_{\mathscr{r}}(A-LC)=\sigma_{\mathscr{r}}(A_{o}-L_{o}e_{n}^{\mathsf{T}})$, and this equals the union of the similarity classes of the right zeros of $a_{d}$. ∎
@@
which equals (26). The resulting observer error dynamics matrix $A_{\rm obs}=A-LC$ is identical to (27) and its right spectrum is $\sigma_{\mathscr{r}}(A_{\rm obs})=\{-1,-2\}$ as desired.
@@ IX SIMULATION EXAMPLE
Consider the SISO continuous-time quaternionic system (2) with $A$ and $C$ from (9), $B=\begin{bmatrix}1&\!\mathbf{j}\end{bmatrix}^{\!\mathsf{T}}$, and D=0D\!=\!0. Using the designs in Examples 3 and 4, we obtain the output-injection gain $L$ (see (26), (32)) that places the observer-error poles at $\lambda_{1}=-1$ and $\lambda_{2}=-2$.
With a step input $u(t)=1-\mathbf{i}+2\mathbf{j}-2\mathbf{k}$, initial states $x_{0}=\bigl[-1+\mathbf{i}-2\mathbf{j}+3\mathbf{k},\;1+2\mathbf{i}-\mathbf{j}-2\mathbf{k}\bigr]^{\mathsf{T}}$, and $\hat{x}_{0}=\bigl[0,0\bigr]^{\mathsf{T}}$, the simulation shows $\hat{x}(t)$ converging to $x(t)$ in all quaternion components, consistent with two stable real poles; see Fig. 1.
@@ X CONCLUSIONS
We developed a characteristic-polynomial-free observer framework for quaternionic systems, centered on an observable companion realization that handles general quaternionic targets, and a quaternionic Ackermann formula restricted to real-coefficient targets. The method enables direct placement of right-eigenvalue similarity classes in the error dynamics and clarifies controller–observer duality over $\mathbb{H},$ with the companion-coordinate coefficient update accommodating noncentral coefficients, whereas Ackermann does not. Future work will extend the approach to multi-input and discrete-time settings.

### v3

@@ Abstract
We develop observer design over hypercomplex quaternions in a characteristic-polynomial-free framework. Using the standard right-module convention, we derive a right observable companion form and companion polynomial that encode error dynamics through right-eigenvalue similarity classes. We also give an Ackermann-type formula for real-coefficient target polynomials, where polynomial evaluation is similarity-equivariant. The resulting recipes place observer poles directly over quaternions and clarify when companion-coordinate updates and one-shot Ackermann formulas remain valid.
@@ I INTRODUCTION
In this paper, we develop an observer-focused, determinant–free framework built on observability and its duality with controllability. We construct the observable companion (observability-form) realization directly from the observability matrix, link its right spectrum to companion-polynomial zeros, and derive a quaternionic Observer Ackermann formula that assigns right–eigenvalue similarity classes using a desired real-coefficient polynomial. Throughout, we remain in the hypercomplex (scalar) model of $\mathbb{H}$ where signals, states, and operators live directly in $\mathbb{H}$. We work with the right spectrum and use left substitution for the companion polynomial—choices that make noncommutativity explicit and usable.
From the application viewpoint, quaternion-based models appear in several areas, but are often handled after conversion to real $4$-vector coordinates, especially in aerospace and robotics, where effective real-vector observers and filters are standard [13]. The purpose of the present paper is different: we show that observer design can also be developed directly over $\mathbb{H}$ in a way that preserves the native hypercomplex structure. This matters precisely in settings where one wants to reason in terms of quaternionic spectra, similarity classes, companion forms, and central polynomial evaluation rather than through coordinate embeddings.
Such a viewpoint is natural not only in quaternionic system theory itself, but also in areas where the hypercomplex structure is intrinsic or algorithmically meaningful, including quantum dynamics and control [5, 12], NMR/MRI pulse design [6], polarization/bivariate signal processing and color imaging [10], quaternion-valued filtering [22], quantum computing [14], and emerging quaternionic learning models [21, 15]. Application-oriented quaternionic observer constructions for four-channel systems are discussed in [20].
We believe the observer results established here over $\mathbb{H}$—including the observable companion construction and the quaternionic Observer Ackermann formula—are original.
@@ III QUATERNIONIC STATE–SPACE SISO SYSTEMS
With $u\!\equiv\!0$, stability means $x(t)\!\to\!0$ (CT) or $x(k)\!\to\!0$ (DT) for all initial states. Equivalently, using the standard right eigenvalues $\lambda$ of $A$, the system is stable iff $\operatorname{Re}\lambda<0$ for all $\lambda$; see [16].
@@ IV OBSERVABLE COMPANION FORM AND COMPANION POLYNOMIAL
    Hence, the right spectrum of $A_{o}$ consists exactly of the similarity classes of the right zeros of $a(\lambda)$.
@@
    Hence, the right spectrum of $A$ consists exactly of the similarity classes of the right zeros of $a(\lambda)$.
@@
When $A$ is cyclic, it is natural to speak of the companion polynomial of $A$ constructed from any observable SISO pair $(A,C)$. Different cyclic choices of $C$ may yield different polynomials. Still, their right zeros agree up to quaternionic similarity classes: for every such $C$, the right zeros of the associated polynomial determine the same set of similarity classes, namely the right spectrum of $A$. For a fixed pair $(A,C)$, however, the observable companion $(A_{o},C_{o})$ and the attached polynomial are unique.
@@ V OBSERVERS
An observer is a dynamic estimator that reconstructs the state $\hat{x}$ from known inputs $u$ and measured outputs $y$ by injecting the innovation $y-C\hat{x}$ through a gain $L$. It enables implementable state feedback $u=K\hat{x}$, monitoring, and prediction. For quaternionic systems, the same observer structure is used, but pole assignment must be interpreted in terms of right-eigenvalue similarity classes.
@@ VI DUALITY OVER $\mathbb{H}$
Over $\mathbb{R}$, controller–observer duality is expressed through transpose: controllability of $(A,B)$ is equivalent to observability of $(A^{\mathsf{T}},B^{\mathsf{T}})$, and one typically transfers gains via $L=K^{\mathsf{T}}$; see, e.g., [2].
Over $\mathbb{H}$, the transpose is no longer the correct dual object. The appropriate operation is the conjugate transpose $M^{*}=\overline{M}^{\mathsf{T}}$, since it preserves the quaternionic Hermitian form, $\langle Ax,y\rangle=\langle x,A^{*}y\rangle,$ and is compatible with quaternionic spectral theory [23, 18, 7, 1]. Thus the duality mapping must be written as
$A\leftrightarrow A^{*},\qquad B\leftrightarrow C^{*},\qquad C\leftrightarrow B^{*},\qquad D\leftrightarrow D^{*}.$
Accordingly, the original system and its dual are linked by
@@
because $(AB)^{*}=B^{*}A^{*}$; see [23, 18]. Since rank is preserved under , this yields
@@
Therefore, if $K\in\mathbb{H}^{1\times n}$ is a state-feedback gain for the controllable dual pair $(A^{*},C^{*})$, then the corresponding observer gain for the original system is
@@
Moreover,
$\sigma_{\mathscr{r}}(A-LC)=\sigma_{\mathscr{r}}(A^{*}-C^{*}K),$
because the adjoint converts right eigenvalues into conjugates, and quaternionic similarity classes satisfy $q]=[\overline{q}]$; see [7, 1]. Hence both matrices have the same assigned right-eigenvalue similarity classes. See also quaternion-valued system treatments in [9, 4].
@@ VII QUATERNIONIC OBSERVER POLE PLACEMENT VIA COMPANION FORM
$A_{\mathrm{obs}}:=A-LC$
consists exactly of the similarity classes induced by the right roots of $a_{d}$, counted with multiplicity. Moreover, with $L_{o}:=T^{-1}L$ we have
@@
has right spectrum $\sigma_{\mathscr{r}}(A_{\mathrm{obs}})=\{[-1],[-2]\}$, with both classes singleton.
@@
Here the assigned class is $\left[-1+\mathbf{i}\right]\in\sigma_{\mathscr{r}}(A_{\mathrm{obs}})$; equivalently, its standard representatives are $1\pm\mathbf{i}$.
(c) Quaternionic coefficients: Choosing target similarity classes represented, for example, by $\lambda_{1}=-1+\mathbf{j}$ and $\lambda_{2}=-2+\mathbf{k},$ sets (rounded)
@@
Its right spectrum consists of the two similarity classes
$\sigma_{\mathscr{r}}(A_{\mathrm{obs}})=\{[-1+\mathbf{i}],[-2+\mathbf{i}]\}=\{[-1+\mathbf{j}],[-2+\mathbf{k}]\}.$
Thus the observable-form coefficient update still assigns the desired similarity classes even when the coefficients are quaternionic.
@@ VIII OBSERVER GAIN VIA ACKERMANN’S FORMULA
Having established the quaternionic observer in its canonical (observable) form, we now turn to a compact algebraic method for directly computing the observer gain. The resulting Observer Ackermann’s formula is the dual counterpart of the control-side expression. It provides a one–shot computation of $L$ achieving any admissible placement of right–eigenvalue classes (observer poles). Its structure mirrors the classical real/complex case. Yet, its justification over $\mathbb{H}$ requires special care: the argument avoids determinants and adjugates and instead exploits the observable companion relations and the right–polynomial calculus developed earlier. This formulation establishes the quaternionic duality between controllability and observability in a determinant–free setting.
@@
Let $A\in\mathbb{H}^{n\times n}$, $C\in\mathbb{H}^{1\times n}$ be such that the SISO pair $(A,C)$ is observable and the observability matrix $\mathcal{O}$ from (3) is invertible. Let
$a_{d}(\lambda)=d_{0}+d_{1}\lambda+\cdots+d_{n-1}\lambda^{n-1}+\lambda^{n}\in\mathbb{R}[\lambda]$
be the desired monic error-dynamics polynomial with real coefficients, and let $e_{n}=[0,0,\dots,1]^{\mathsf{T}}\in\mathbb{R}^{n}$. Then the observer gain column
@@
$\sigma_{\mathscr{r}}(A-LC)$ $=$ $\{[\lambda_{i}]:i=1,\ldots,n\},$ (28)
where $\{\lambda_{i}\}_{i=1}^{n}$ are right roots of $a_{d}$ (listed with multiplicity).
@@
The restriction
$a_{d}\in\mathbb{R}[\lambda]$
is essential: real coefficients lie in the center $Z(\mathbb{H})=\mathbb{R}$, so $a_{d}(A)$ is unambiguous and the required polynomial–similarity interchange is valid. The assigned classes are real or complex-conjugate spherical: for an $n$-th order SISO design, each real class consumes one degree and each nonreal class two, so $2s+r\leq n$ when placing $s$ nonreal and $r$ real classes.
@@
Similarity preserves the right spectrum, hence $\sigma_{\mathscr{r}}(A-LC)=\sigma_{\mathscr{r}}(A_{o}-L_{o}e_{n}^{\mathsf{T}})$, and this spectrum consists exactly of the similarity classes of the right zeros of $a_{d}$. ∎
@@
which equals (25). The resulting observer error dynamics matrix $A_{\rm obs}=A-LC$ is identical to (26) and its right spectrum is $\sigma_{\mathscr{r}}(A_{\mathrm{obs}})=\{[-1],[-2]\}$, as desired, with both classes singleton.
@@ IX SIMULATION EXAMPLE
For (2) with $A,C$ from (9), $B=\begin{bmatrix}1&\!\mathbf{j}\end{bmatrix}^{\!\mathsf{T}}$, $D=0$, input $u(t)=1-\mathbf{i}+2\mathbf{j}-2\mathbf{k}$, and initial conditions $x_{0}=\bigl[-1+\mathbf{i}-2\mathbf{j}+3\mathbf{k},\;1+2\mathbf{i}-\mathbf{j}-2\mathbf{k}\bigr]^{\mathsf{T}}$, $\hat{x}_{0}=\bigl[0,0\bigr]^{\mathsf{T}}$, the gain from (25)–(31) yields convergence $\hat{x}(t)\to x(t)$ in all quaternionic components; see Fig. 1.
@@ X CONCLUSIONS
We developed a characteristic-polynomial-free observer framework for quaternionic systems, centered on an observable companion realization and a quaternionic Ackermann formula for real-coefficient targets. The method assigns right-eigenvalue similarity classes directly in the error dynamics and makes the -based controller–observer duality explicit; companion-coordinate updates accommodate noncentral coefficients, whereas Ackermann does not. Application-oriented quaternionic observers for four-channel systems are discussed in [20], while related controllability companion-form and state-feedback results are developed in [19].